# Quasi-boolean groups

### v2

@@ 5. Maximal pro-$2$ quotients of real projective groups
Proof.
We need the following group-theoretical
@@ 2. $(ii)$
Let $2^{b}$ denote the order of $D$. Let $N$ be the kernel of $f$ and write the order of $N$ as $2^{a}r$ where $r$ is not divisible by $2$. Then the order of $C$ is $2^{a+b}r$, so the order of $P$ is $2^{a+b}$, while the order of $P\cap N$ is at most $2^{a}$. Since the kernel of the restriction $f|_{P}\colon P\to H$ is $P\cap N$, we get that the image of $f|_{P}$ is at least $2^{b}$. Therefore $f|_{P}$ is surjective.
Since the order of the subgroup generated by $g$ divides $2$ there is a $t\in C$ such that $t^{-1}gt\in P$ by the second Sylow theorem. By the above there is a $v\in P$ such that $f(v)=f(t)$. Set $h=tv^{-1}$. Then
$h^{-1}gh=v(t^{-1}gt)v^{-1}\in vPv^{-1}=P,$
since $v^{-1}\in P$. Moreover
$f(h^{-1}gh)=f(v)f(t)^{-1}f(g)f(t)f(v)^{-1}=f(g)$
using that $f$ is a homomorphism and $f(v)=f(t)$. ∎
@@
be a real embedding problem $\mathbf{E}$ for $G$. Let $H\subseteq A$ be the image of $\phi$. Since $G$ is a pro-$2$ group, $H$ is a $2$-group. Let $C\subseteq B$ be the pre-image of $H$ with respect to $\alpha$ and let $P\subseteq C$ be a $2$-Sylow subgroup. By part $(i)$ of Lemma 5.3 the restriction $\alpha|_{P}$ is surjective. Clearly,
@@
is a $2$-embedding problem $\mathbf{F}$ for $G$ such that if it has a solution then $\mathbf{E}$ also has a solution. Therefore it will be enough show that $\mathbf{F}$ is real because of our assumptions on $G$. Let $x\in G$ be an involution; by assumption there is a $2$-torsion element $g\in C$ such that $\alpha(g)=\phi(x)$. Then there is a $y\in P$ which is conjugate to $g$ in $C$ such that $\alpha(y)=\phi(x)$ by part $(ii)$ of Lemma 5.3. Since $y$ is conjugate to a $2$-torsion element, it is also $2$-torsion. So $\mathbf{F}$ is real. ∎
@@
has a solution. Now let $K$ be a pseudo real closed field such that $G$ is isomorphic to the maximal pro-$2$ quotient of the absolute Galois group $\Gamma$ of $K$. Such a field exists by the work of Haran–Jarden in [5, 6]. Let $q\colon\Gamma\to G$ be the corresponding quotient homomorphism. We claim that
@@
is a real embedding problem $\mathbf{F}$. Indeed, let $x\in\Gamma$ be an involution such that $\phi\circ q(x)$ is also an involution. Then $q(x)\in G$ is also an involution, so there is an involution $g\in B$ such that $\alpha(g)=\phi(q(x))=\phi\circ q(x)$. So $\mathbf{F}$ is real. Since $\Gamma$ is real projective, the embedding problem $\mathbf{F}$ has a solution $\widetilde{\phi}\colon\Gamma\to B$. But $B$ is a $2$-group, so $\widetilde{\phi}$ is the composition of $q$ and a continuous homomorphism $G\to B$. The latter is a solution to $\mathbf{E}$. To finish the proof of Theorem 5.4 we need the following notation and lemma.
@@
Let $G$ be a profinite group, let $H\subseteq G_{2}$ be an open subgroup, let $I$ be the pre-image $t_{G}^{-1}(H)\subseteq G$, and let $h\colon I\to H$ denote the restriction of $t_{G}$ onto $I$. Then $h_{2}\colon I_{2}\to H_{2}=H$ is an isomorphism.
@@
Because $h$ is surjective, the map $h_{2}$ is also surjective, so we only need to show that it is injective, too. Let $1\neq\gamma\in I_{2}$ be arbitrary. Then there is an open, normal subgroup $U\subseteq I_{2}$ such that $\gamma\not\in U$. Since $t_{I}$ is continuous, the pre-image $t_{I}^{-1}(U)$ is an open subgroup of $2$-power index in $I$. Since $I$ is an open subgroup of $2$-power index in $G$, we get that $t_{I}^{-1}(U)$ is an open subgroup of $2$-power index in $G$, too. Set $N=\bigcap_{\delta\in G}\delta^{-1}t_{I}^{-1}(U)\delta$; clearly it is a normal subgroup. Since $\delta^{-1}t_{I}^{-1}(U)\delta$ only depends on the coset $t_{I}^{-1}(U)\delta$, of which there are only finitely many, we get that $N$ is a finite intersection of open subgroups of $2$-power index in $G$, so it is also an open subgroup of $2$-power index in $G$. Therefore the subgroup $t_{I}(N)\subseteq I_{2}$ is the pre-image of $t_{G}(N)\subseteq G_{2}$, but clearly $\gamma\not\in t_{I}(N)$, and hence $h_{2}(\gamma)\neq 1$. ∎
Now we return to the proof of Theorem 5.4. By Definition 1.1, it remains to show that $G=\Gamma_{2}$ contains an open subgroup without $2$-torsion. Let $K$ be the field introduced above, and let $\Delta\subseteq\Gamma$ be the open subgroup corresponding to the finite extension $K(\sqrt{-1})/K$. As $\Delta$ is isomorphic to the absolute Galois group of $K(\sqrt{-1})$, it has cohomological dimension at most $1$. By the Rost–Voevodsky norm residue theorem [14] (formerly known as the Milnor conjecture), the pull-back map $H^{\bullet}(\Delta_{2},\mathbb{Z}/2)\to H^{\bullet}(\Delta,\mathbb{Z}/2)$ induced by the quotient map $t_{\Delta}\colon\Delta\to\Delta_{2}$ is an isomorphism. Therefore, $\Delta_{2}$ also has cohomological dimension at most $1$. Hence, by [12, Proposition 14 on page 19], any closed subgroup of $\Delta_{2}$ has cohomological dimension at most $1$. Since a finite subgroup would have infinite cohomological dimension, $\Delta_{2}$ is torsion-free. Moreover, $\Delta$ has index dividing two in $\Gamma$, so $\Delta_{2}$ is isomorphic to the kernel of the homomorphism $\Gamma_{2}\to\mathbb{Z}/2\mathbb{Z}$ corresponding to the homomorphism $\Gamma\to\mathrm{Gal}(K(\sqrt{-1})/K)\subseteq\mathbb{Z}/2\mathbb{Z}$ by Lemma 5.6. Therefore, $\Delta_{2}$ is an open, torsion-free subgroup of $\Gamma_{2}$. This finishes the proof of Theorem 5.4. ∎
@@ 6. Pro-$2$ real projective groups versus quasi-Boolean groups
The goal of this section is to prove Theorems 6.6 and 6.17. We begin with the following recollection and notation.
@@
The free product of pro-$p$ groups $G_{1}$ and $G_{2}$ is the coproduct of $G_{1}$ and $G_{2}$ in the category of pro-$p$ groups, that is, it has the following universal property. For $j=1,2$, let $\iota_{j}\colon G_{j}\to G_{1}*_{p}G_{2}$ be the composition of the natural inclusion $G_{j}\to G_{1}*G_{2}$ and the quotient homomorphism $G_{1}*G_{2}\to G_{1}*_{p}G_{2}$. Then for every pro-$2$ group $G$ and for every pair of homomorphisms $f_{j}\colon G_{j}\to G$ of pro-$2$ groups, there is a unique homomorphism $f_{1}*_{p}f_{2}\colon G_{1}*_{p}G_{2}\to G$ of pro-$2$ groups such that $(f_{1}*_{p}f_{2})\circ\iota_{j}=f_{j}$ for $j=1,2$. This follows obviously from the definition when $G$ is finite, and the general case follows by taking the projective limit.
Definition 6.2.
Let $G$ be a profinite group. Let $\mathcal{Y}_{p}(G)$ denote the subset of elements of order dividing $p$ in $G$. We equip $\mathcal{Y}_{p}(G)$ with the subset topology. Let $\mathcal{X}_{p}(G)$ denote the quotient of $\mathcal{Y}_{p}(G)$ by the conjugation action of $G$. We equip $\mathcal{X}_{p}(G)$ with the quotient topology. Let $\mathcal{Y}_{p}^{*}(G)$ denote the complement of $1$ in $\mathcal{Y}_{p}(G)$, and let $\mathcal{X}_{p}^{*}(G)$ denote the complement of the conjugacy class of $1$ in $\mathcal{X}_{p}(G)$. When $p=2$ we let $\mathcal{Y}(G),\mathcal{X}(G),\mathcal{Y}^{*}(G),\mathcal{X}^{*}(G)$ denote $\mathcal{Y}_{p}(G),\mathcal{X}_{p}(G),\mathcal{Y}_{p}^{*}(G),\mathcal{X}_{p}^{*}(G)$, respectively.
@@
Let $X$ be a topological space. Recall from Definition 1.4 the free pro-$2$ power $\mathbb{B}(X)$. We note that $\mathbb{B}(X)$ has the following universal property: Let $\iota_{X}\colon X\to\mathbb{B}(X)$ be the composition of the natural inclusion $X\to\bigast_{X}\mathbb{Z}/2\mathbb{Z}$ and the quotient homomorphism $\bigast_{X}\mathbb{Z}/2\mathbb{Z}\to\mathbb{B}(X)$. Then, for every pro-$2$ group $G$ and for every continuous map $f\colon X\to\mathcal{Y}(G)$, there is a unique homomorphism $b_{f}\colon\mathbb{B}(X)\to G$ of pro-$2$ groups such that $b_{f}\circ\iota_{X}=f$. This follows obviously from the definition when $G$ is finite, and the general case follows by taking the projective limit.
@@
Let $f\colon X\to Y$ be a continuous map of topological spaces. Since $i_{Y}\colon Y\to\mathbb{B}(Y)$ is continuous and its image lies in $\mathcal{Y}(\mathbb{B}(Y))$, by the universal property in the remark above there is a unique homomorphism $\mathbb{B}(f)\colon\mathbb{B}(X)\to\mathbb{B}(Y)$ of pro-$2$ groups such that $\mathbb{B}(f)\circ\iota_{X}=\iota_{Y}\circ f$. This makes the assignment $X\mapsto\mathbb{B}(X)$ into a functor.
@@
For every topological space $X$, the map $\mathbb{B}(u_{X})\colon\mathbb{B}(X)\to\mathbb{B}(\widehat{X})$ induced by the profinite completion $u_{X}\colon X\to\widehat{X}$ is an isomorphism of pro-$2$ groups.
@@
First we are going to show that $\mathbb{B}(u_{X})$ is surjective. Since the image is compact, it is closed, so it will be sufficient to show that the image is dense. In order to do so it will be enough to prove that, for every $2$-group $G$ and continuous surjective homomorphism $f\colon\mathbb{B}(\widehat{X})\to G$, the composition $f\circ\mathbb{B}(u_{X})$ is surjective. Note that $\widehat{X}$ generates $\bigast_{\widehat{X}}\mathbb{Z}/2\mathbb{Z}$, so $f(\widehat{X})$ generates $G$. Since $u_{X}(X)$ is dense in $\widehat{X}$, the image $f\circ u_{X}(X)$ is dense in $f(\widehat{X})$. But $G$ is finite, so it is discrete, and hence $f\circ u_{X}(X)$ is equal to $f(\widehat{X})$. So $f\circ u_{X}(X)$ generates $G$, therefore $f\circ\mathbb{B}(u_{X})$ is surjective.
Next we are going to show that $\mathbb{B}(u_{X})$ is injective. In order to do so it will be sufficient to show that $u_{X}\circ\iota_{X}\colon X\to\mathbb{B}(\widehat{X})$ has the universal property in Remark 6.3. Indeed then there is a continuous homomorphism $f\colon\mathbb{B}(\widehat{X})\to\mathbb{B}(X)$ such that $f\circ\mathbb{B}(u_{X})\circ\iota_{X}$ is $\iota_{X}$, and hence $f\circ\mathbb{B}(u_{X})$ is the identity of $\mathbb{B}(X)$. Now let $G$ be a pro-$2$ group and $f\colon X\to\mathcal{Y}(G)$ be a continuous map. Since $\mathcal{Y}(G)$ is profinite, there is a unique continuous map $g\colon\widehat{X}\to\mathcal{Y}(G)$ such that $f=g\circ u_{X}$ because of the universal property of $u_{X}$. Using the universal property of $\mathbb{B}(\widehat{X})$ we get that there is a continuous homomorphism $b_{g}\colon\mathbb{B}(\widehat{X})\to G$ of pro-$2$ groups such that $b_{f}\circ\iota_{\widehat{X}}=g$. ∎
@@
Let $G$ be a quasi-Boolean pro-$2$ group. It will be sufficient to show that $G$ is the maximal pro-$2$ quotient of a real projective group by Theorem 5.4. In order to do so, we will use a group-theoretical characterisation of real projective groups by Haran and Jarden. Following [6, Definition 1.1 on page 156] we define:
@@
This claim is [6, Theorem 3.6 on page 160]. ∎
Now we return to the proof of Theorem 6.6. For every set $Y$, let $F(Y)$ denote the free pro-$2$ group on $Y$ as defined in [12, Section 1.5 on pages 7–8]. By assumption, $G$ is the free product of a free pro-$2$ group $F(Y)$ and a Boolean group $\mathbb{B}(X)$ for a set $Y$ and a profinite space $X$. By Theorem 6.8 it will be sufficient to show that $G$ is the maximal pro-$2$ quotient of a real free profinite group. Let $\widehat{G}$ be the free product $\bigast_{X}\mathbb{Z}*\bigast_{X}\mathbb{Z}/2\mathbb{Z}$, i.e., the group which is freely generated by the elements of the disjoint union of $Y$ and $X$, subject to the relation that these elements are involutions. Let $\mathcal{N}$ be the family of normal subgroups $N$ of $\widehat{G}$ such that

### v3

@@ 5. Maximal pro-$2$ quotients of real projective groups
To prove Proposition 5.2 we need the following group-theoretical
@@ 2. $(ii)$
Let $2^{b}$ denote the order of $D$. Let $N$ be the kernel of $f$ and write the order of $N$ as $2^{a}r$ where $r$ is not divisible by $2$. Then the order of $C$ is $2^{a+b}r$, so the order of $P$ is $2^{a+b}$, while the order of $P\cap N$ is at most $2^{a}$. Since the kernel of the restriction $f|_{P}\colon P\to D$ is $P\cap N$, we get that the image of $f|_{P}$ is at least $2^{b}$. Therefore $f|_{P}$ is surjective. This proves $(i)$.
Since the order of the subgroup generated by $x$ divides $2$, there is a $t\in C$ such that $t^{-1}xt\in P$ by the second Sylow theorem. By part $(i)$, there is a $v\in P$ such that $f(v)=f(t)$. Set $h=tv^{-1}$. Then
$h^{-1}xh=v(t^{-1}xt)v^{-1}\in vPv^{-1}=P,$
since $v^{-1}\in P$. Moreover,
$f(h^{-1}xh)=f(v)f(t)^{-1}f(x)f(t)f(v)^{-1}=f(x)$
using that $f$ is a homomorphism and $f(v)=f(t)$. This proves $(ii)$. ∎
Proof of Proposition 5.2.
@@
be a real embedding problem $\mathbf{E}$ for $G$. Let $H\subseteq A$ be the image of $\phi$. Since $G$ is a pro-$2$ group, $H$ is a $2$-group. Let $C\subseteq B$ be the preimage of $H$ with respect to $\alpha$ and let $P\subseteq C$ be a $2$-Sylow subgroup. By part $(i)$ of Lemma 5.3 the restriction $\alpha|_{P}\colon P\to H$ is surjective. Clearly,
@@
is a $2$-embedding problem $\mathbf{F}$ for $G$ such that if it has a solution then $\mathbf{E}$ also has a solution. Therefore it will be enough to show that $\mathbf{F}$ is real because of our assumptions on $G$. Let $x\in G$ be an involution. By assumption there is a $2$-torsion element $g\in C$ such that $\alpha(g)=\phi(x)$. Then there is a $y\in P$ which is conjugate to $g$ in $C$ such that $\alpha(y)=\phi(x)$ by part $(ii)$ of Lemma 5.3. Since $y$ is conjugate to a $2$-torsion element, it is also $2$-torsion. So $\mathbf{F}$ is real. ∎
@@
has a solution. Recall that a field $K$ is called pseudo real closed if every absolutely irreducible variety defined over $K$ which has a simple $K_{<}$-rational point for every ordering << on $K$ has a $K$-rational point, where $K_{<}$ denotes the real closure of the ordered field $(K,<)$. By the assumption, there is a real projective group $\Gamma$ such that $G$ is isomorphic to the maximal pro-$2$ quotient of $\Gamma$. By the work of Haran–Jarden in [9, Theorem 10.4 on page 487], we can then choose a pseudo real closed field $K$ such that $\Gamma$ is the absolute Galois group of $K$. Let $q\colon\Gamma\to G$ be the corresponding quotient homomorphism. We claim that
@@
is a real embedding problem which we denote by $\mathbf{F}$. Indeed, let $x\in\Gamma$ be an involution such that $\phi\circ q(x)$ is also an involution. Then $q(x)\in G$ is also an involution, so there is an involution $g\in B$ such that $\alpha(g)=\phi(q(x))=(\phi\circ q)(x)$. So $\mathbf{F}$ is real. Since $\Gamma$ is real projective as the absolute Galois group of a pseudo real closed field, the embedding problem $\mathbf{F}$ has a solution $\widetilde{\phi}\colon\Gamma\to B$. But $B$ is a $2$-group, so $\widetilde{\phi}$ is the composition of $q$ and a continuous homomorphism $G\to B$. The latter is a solution to $\mathbf{E}$. To finish the proof of Theorem 5.4 we need the following notation and lemma.
@@
Let $G$ be a profinite group, let $H\subseteq G_{2}$ be an open subgroup, let $I$ be the preimage $t_{G}^{-1}(H)\subseteq G$, and let $h\colon I\to H$ denote the restriction of $t_{G}$ onto $I$. Then $h_{2}\colon I_{2}\to H_{2}=H$ is an isomorphism.
@@
Since $h$ is surjective, the map $h_{2}$ is also surjective, so we only need to show that it is injective, too. Let $1\neq\gamma\in I_{2}$ be arbitrary. Then there is an open normal subgroup $U\subseteq I_{2}$ such that $\gamma\not\in U$. Since $t_{I}$ is continuous, the preimage $t_{I}^{-1}(U)$ is an open subgroup of $2$-power index in $I$. Since $I$ is an open subgroup of $2$-power index in $G$, we get that $t_{I}^{-1}(U)$ is an open subgroup of $2$-power index in $G$, too. Set $N\coloneqq\bigcap_{\delta\in G}\delta^{-1}t_{I}^{-1}(U)\delta$. Clearly $N$ is a normal subgroup. Since $\delta^{-1}t_{I}^{-1}(U)\delta$ only depends on the coset $t_{I}^{-1}(U)\delta$, of which there are only finitely many, we get that $N$ is a finite intersection of open subgroups of $2$-power index in $G$, so it is also an open subgroup of $2$-power index in $G$. Therefore the subgroup $t_{I}(N)\subseteq I_{2}$ is the preimage of $t_{G}(N)\subseteq G_{2}$, but clearly $\gamma\not\in t_{I}(N)$, and hence $h_{2}(\gamma)\neq 1$. ∎
Now we return to the proof of Theorem 5.4. By Definition 1.1, it remains to show that $G=\Gamma_{2}$ contains an open subgroup without $2$-torsion. Let $K$ be the field introduced above, and let $\Delta\subseteq\Gamma$ be the open subgroup corresponding to the finite extension $K(\sqrt{-1})/K$. As $\Delta$ is isomorphic to the absolute Galois group of $K(\sqrt{-1})$, it has cohomological dimension at most $1$ by [15, Corollary 2.4]. By the Rost–Voevodsky norm residue theorem [23] (formerly known as the Milnor conjecture), the pull-back map $H^{\bullet}(\Delta_{2},\mathbb{Z}/2\mathbb{Z})\to H^{\bullet}(\Delta,\mathbb{Z}/2\mathbb{Z})$ induced by the quotient map $t_{\Delta}\colon\Delta\to\Delta_{2}$ is an isomorphism. Therefore, $\Delta_{2}$ also has cohomological dimension at most $1$. Hence, by [21, Proposition 14 on page 19], any closed subgroup of $\Delta_{2}$ has cohomological dimension at most $1$. Since a finite subgroup would have infinite cohomological dimension, $\Delta_{2}$ is torsion-free. Moreover, $\Delta$ has index dividing two in $\Gamma$, so $\Delta_{2}$ is isomorphic to the kernel of the homomorphism $\Gamma_{2}\to\mathbb{Z}/2\mathbb{Z}$ corresponding to the homomorphism $\Gamma\to\mathrm{Gal}(K(\sqrt{-1})/K)\subseteq\mathbb{Z}/2\mathbb{Z}$ by Lemma 5.6. Therefore, $\Delta_{2}$ is an open, torsion-free subgroup of $\Gamma_{2}$. This finishes the proof of Theorem 5.4. ∎
@@ 6. Pro-$2$ real projective groups versus quasi-Boolean groups
The goal of this section is to prove Theorems 6.6 and 6.18. We begin with the following recollection and notation.
@@
The free product $G_{1}*_{p}G_{2}$ of pro-$p$ groups $G_{1}$ and $G_{2}$ is the coproduct of $G_{1}$ and $G_{2}$ in the category of pro-$p$ groups, that is, it has the following universal property. For $j=1,2$, let $\iota_{j}\colon G_{j}\to G_{1}*_{p}G_{2}$ be the composition of the natural inclusion $G_{j}\to G_{1}*G_{2}$ and the quotient homomorphism $G_{1}*G_{2}\to G_{1}*_{p}G_{2}$. Then for every pro-$2$ group $G$ and for every pair of homomorphisms $f_{j}\colon G_{j}\to G$ of pro-$2$ groups, there is a unique homomorphism $f_{1}*_{p}f_{2}\colon G_{1}*_{p}G_{2}\to G$ of pro-$2$ groups such that $(f_{1}*_{p}f_{2})\circ\iota_{j}=f_{j}$ for $j=1,2$. This follows obviously from the definition when $G$ is finite, and the general case follows by taking the projective limit.
Notation 6.2.
Let $G$ be a profinite group. Let $\mathcal{Y}_{p}(G)$ denote the subset of elements of order dividing $p$ in $G$. We equip $\mathcal{Y}_{p}(G)$ with the subspace topology. Let $\mathcal{X}_{p}(G)$ denote the quotient of $\mathcal{Y}_{p}(G)$ by the conjugation action of $G$. We equip $\mathcal{X}_{p}(G)$ with the quotient topology. Let $\mathcal{Y}_{p}^{*}(G)$ denote the complement of $1$ in $\mathcal{Y}_{p}(G)$, and let $\mathcal{X}_{p}^{*}(G)$ denote the complement of the conjugacy class of $1$ in $\mathcal{X}_{p}(G)$. When $p=2$ we let $\mathcal{Y}(G),\mathcal{X}(G),\mathcal{Y}^{*}(G),\mathcal{X}^{*}(G)$ denote $\mathcal{Y}_{p}(G),\mathcal{X}_{p}(G),\mathcal{Y}_{p}^{*}(G),\mathcal{X}_{p}^{*}(G)$, respectively. We note that $\mathcal{Y}^{*}(G)$ equals the set $\mathrm{Inv}(G)$ of involutions in $G$.
@@
Let $X$ be a topological space. Recall from Definition 1.5 the free pro-$2$ product $\mathbb{B}(X)$. We note that $\mathbb{B}(X)$ has the following universal property: Let $\iota_{X}\colon X\to\mathbb{B}(X)$ be the composition of the natural inclusion $X\to\bigast_{X}\mathbb{Z}/2\mathbb{Z}$ and the quotient homomorphism $\bigast_{X}\mathbb{Z}/2\mathbb{Z}\to\mathbb{B}(X)$. Then, for every pro-$2$ group $G$ and for every continuous map $f\colon X\to\mathcal{Y}(G)$, there is a unique homomorphism $b_{f}\colon\mathbb{B}(X)\to G$ of pro-$2$ groups such that $b_{f}\circ\iota_{X}=f$. This follows from the definition when $G$ is finite, and the general case follows by taking the projective limit.
@@
Let $f\colon X\to Y$ be a continuous map of topological spaces. Since $i_{Y}\colon Y\to\mathbb{B}(Y)$ is continuous and its image lies in $\mathcal{Y}(\mathbb{B}(Y))$, by the universal property in Remark 6.3 above there is a unique homomorphism $\mathbb{B}(f)\colon\mathbb{B}(X)\to\mathbb{B}(Y)$ of pro-$2$ groups such that $\mathbb{B}(f)\circ\iota_{X}=\iota_{Y}\circ f$. This makes the assignment $X\mapsto\mathbb{B}(X)$ into a functor.
@@
For every topological space $X$, the map $𝔹(X^)\mathbb{B}(u_{X})\colon\mathbb$$\to\mathbb{B}(\widehat{X})$ induced by the profinite completion $u_{X}\colon X\to\widehat{X}$ is an isomorphism of pro-$2$ groups.
@@
[…]
$b_{f}\circ\mathbb{B}(u_{X})\circ\iota_{X}=b_{f}\circ\iota_{\widehat{X}}\circ u_{X}=g\circ u_{X}=f,$
as desired. ∎
@@
Let $G$ be a quasi-Boolean pro-$2$ group. It will be sufficient to show that $G$ is the maximal pro-$2$ quotient of a real projective group by Theorem 5.4. In order to do so, we will use a group-theoretical characterisation of real projective groups by Haran and Jarden. Following [10, Definition 1.1 on page 156] we define:
@@
This claim is [10, Theorem 3.6 on page 160]. ∎
Now we return to the proof of Theorem 6.6. For every set $Y$, let $F(Y)$ denote the free pro-$2$ group on $Y$ as defined in [21, Section 1.5 on pages 7–8]. By assumption, $G$ is the free product of a free pro-$2$ group $F(Y)$ and a Boolean group $\mathbb{B}(X)$ for a set $Y$ and a profinite space $X$. By Theorem 6.8 it will be sufficient to show that $G$ is the maximal pro-$2$ quotient of a real free profinite group. Let $\widehat{G}$ be the free product $\bigast_{Y}\mathbb{Z}*\bigast_{X}\mathbb{Z}/2\mathbb{Z}$, i.e., the group which is freely generated by the elements of the disjoint union of $Y$ and $X$, subject to the relation that the elements in $X$ are involutions. Let $\mathcal{N}$ be the family of normal subgroups $N$ of $\widehat{G}$ such that